# Diffusion2: Dynamic 3D Content Generation via Score Composition of Orthogonal Diffusion Models

### v3

@@ 3.1 Image matrix generation
$=\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)+\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)-\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right).\qed$
Here $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)$ and $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)$ are just the score functions of the video diffusion model and the multi-view diffusion model respectively. Since the $\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right)$ is intractable in the conditional generation, we use the convex combination of $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)$ and $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)$ to replace it as:
$\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right)=(1-s)\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)+s\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right).$ (10)
In practice, we employ a logistic schedule to adjust $s$ at each denoising step. Given the current denoising step $i$ and the number of total steps $N$, we set $s=1-\frac{1}{1+e^{k(i/N-s_{0})}}$. This function has a sigmoidal curve, which is relatively flat at the extremes away from middle $s_{0}$ and changes sharply near it, with the derivative controlled by $k$. This schedule decouples the generation of consistent geometry and temporally smooth appearance to some extent.
Sampling in latent space   For convenience, the above theorem assumes that we sample images in RGB space. However, modern high-resolution diffusion models typically generate images in a latent space encoded by pretrained auto-encoder kingma2014auto ; van2017neural . The legitimacy of the aforementioned derivation requires that the multi-view generator and the video generator share the same latent space. This requirement is not met by most existing instantiation of them. However, as pointed out by SVD blattmann2023stable , video generation models trained on large-scale video datasets have learned a strong stereo knowledge, thus providing a better pretraining for multi-view diffusion models than those trained solely on image data. Therefore, we believe that the requirement will be increasingly satisfied in the future.
Generation with various input conditions   Note that the formulation described above is based on unconditional generation. However, we are more interested in controllable generation in practice. Then we will discuss on how to extend the above process to conditional generation. Consider the augmented matrix $\mathcal{I}_{\text{aug}}$ defined by
Compared to the aforementioned unconditional generation, the augmented matrix $\mathcal{I}_{\text{aug}}$ additionally includes the optional input image $I_{0,0}$, and $\mathcal{I}_{0,\{1:V\}},\mathcal{I}_{\{1:F\},0}$ that represent the reference geometry and dynamic of target object. To incorporate them into the generation of $\mathcal{I}$, we condition the score estimator for denoising each row (or column) in equations (5) and (10) on the first element of the same row (or column) in $\mathcal{I}_{\text{aug}}$. This formulation enables us to handle various 4D generation tasks depending on different forms of conditions specified by users as below:
  • Image-to-4D. Given $I_{0,0}$ as input, both $\mathcal{I}_{0,\{1:V\}}$ and $\mathcal{I}_{\{1:F\},0}$ are generated by diffusion models.
  • Video-to-4D. Given $\mathcal{I}_{\{0:F\},0}$ as input, we use the last frame $I_{0,0}$ as the condition of multi-view diffusion model to generate $\mathcal{I}_{0,\{1:V\}}$.
  • 3D-to-4D. Similarly, Given $\mathcal{I}_{0,\{0:V\}}$ as input, we use the front view $I_{0,0}$ as the condition of video diffusion model to generate $\mathcal{I}_{\{1:F\},0}$.
  • Text-to-4D. The above inputs including single image $I_{0,0}$, single-view video $\mathcal{I}_{\{0:F\},0}$ and static multi-view images $\mathcal{I}_{0,\{0:V\}}$ can all be generated guided by text.
Once $\mathcal{I}_{0,\{1:V\}},\mathcal{I}_{\{1:F\},0}$ are obtained, we can sample the full matrix $\mathcal{I}$ using equations (5) and (10).
Parallel denoising   Assumption 3.1 ensures the safety of independently generating the geometry $\mathcal{I}_{0,\{1:V\}}$ and the motion $\mathcal{I}_{\{1:F\},0}$. Since these two generation processes have no computational or data dependency, their total time cost could be reduced to a single reverse diffusion process. Additionally, when we denoise the rest part of $\mathcal{I}_{\text{aug}}$, i.e., $\mathcal{I}$, the score estimation for each row and column can also be parallelized in each denoising step. Therefore, with sufficient GPU memory, the total time spent on the process illustrated in Figure 2 (ii) remains the same as that for generating a single video.
3.2 Robust reconstruction
4D representation   Given generated synchronized multi-view videos from any type of condition, there are numerous methods that can be employed to reconstruct complete 4D assets. Among these candidates, we adopt the 4D Gaussian Splatting yang2023gs4d due to its superior fitting capabilities and efficient optimization given the dense multi-view supervision.
Optimization   Although the images produced in the first stage already exhibit intuitively satisfactory spatiotemporal consistency, the performance limitation of the foundational diffusion model still makes it difficult to achieve precise pixel-level matching across different views and frames. Therefore, we follow chen2024v3d to optimize the combination of perception loss $\mathcal{L}_{lpips}$ zhang2018unreasonable and D-SSIM $\mathcal{L}_{ssim}$ wang2004image while discarding L1 loss. In addition, we weight each term with the confidence score as pan2024fast , then the total objective is defined as $\mathcal{L}_{total}=\lambda_{lpips}\mathcal{C}_{lpips}\mathcal{L}_{lpips}+\lambda_{ssim}\mathcal{C}_{ssim}\mathcal{L}_{ssim}$, where $\mathcal{C}_{ssim}$ is just the SSIM between ground truth and rendered images and $\mathcal{C}_{lpips}$ is defined as $1-\mathcal{L}_{lpips}$.
4 Experiments
4.1 Implementation details
[…]
(a) Image-to-4D
(b) Video-to-4D
Figure 3: Qualitative comparisons on (a) image-to-4D generation and (b) video-to-4D generation.
4.2 Main results
4D generation from reference image
Table 1: User study on image-to-4D generation. The proportions of different methods that best match user preferences under three criteria are reported.
Method Details Geometrical Temporal Overall model Generation
consistency smoothness quality time
Animate124 zhao2023animate124 11.3% 31.0% 16.0% 18.0% 9h
DreamGaussian4D ren2023dreamgaussian4d 27.7% 24.3% 48.0% 25.7% 12m
Diffusion2 (Ours) 61.0% 44.7% 36.0% 56.3% 10m
In Figure 3 (a), we present the results generated by our method and compare them with other alternatives. It can be observed that our simple and elegant pipeline is capable of generating 4D assets with comparable quality to those produced by SDS-based methods using sophisticated multi-stage optimization. Furthermore, the parallel denoising characteristic can provide an efficiency advantage for our method over those need recurrent evaluation of diffusion UNet. We also conducted a user study in Table 1. It suggests that our method garnered the highest preference in the multi-view consistency, detail, and overall model quality.
4D generation from single view video   Generating 4D dynamic objects from single-view video is a practical task first introduced in jiang2023consistent4d . Compared to image-to-4D generation, this task places additional constraints on the object’s motion. Our proposed framework can be easily adapted to deal with this task as detailed in Section 3.1. We perform both quantitative and qualitative comparisons with other counterparts in this setting. The qualitative result is shown in Figure 3 (b). It can be seen that our method slightly reduces the over-saturated appearance led by SDS while achieving more clear details. For quantitative evaluation, we report the CLIP-similarity between generated views and ground truth images to indicate overall semantic consistency and the recognizability of generated images. As shown in Table 2, our method can achieve better CLIP similarity than other alternatives.
4D generation from static 3D content
Figure 4: Synthesized images from static 3D models.
Table 2: Quantitative comparisons on video-to-4D generation.
Method Type CLIP Similarity $\uparrow$ Generation time $\downarrow$
Consistent4D jiang2023consistent4d Optimization-based 0.87 2h
4DGen yin20234dgen Optimization-based 0.89 2h10m
Efficient4D pan2024fast Photogrammetry-based 0.92 14m
Diffusion2 (Ours) Photogrammetry-based 0.94 10m
Naturally, Diffusion2 can also animate static 3D models as described in Section 3.1, which has substantial practical implications. From Figure 4 we can observe that our method is capable of endowing 3D models with diverse and realistic dynamics while maintaining satisfactory temporal and geometrical consistency. More results including text-to-4D can be found in Section A.5.
4.3 Ablation studies
Figure 5: Ablation studies. (a) & (b) The tuning of parameters $s_{0}$ and $k$ controlling the logistic schedule. (c) Different types of scale schedule. Best viewed with zoom-in.
Since previous methods have explored dynamic content reconstruction well, we mainly focus on the ablation of key design choices in the stage of image matrix generation. Recall that in equation (10), we use $s$ to balance the contribution of two foundation models. And we choose a logistic function $s=1-\frac{1}{1+e^{k(i/N-s_{0})}}$ as the schedule of $s$, which has two hyper-parameters $s_{0}$ and $k$ controlling the shape of the function. Then we tune $s_{0}$ and $k$ to study their impact on generation.
Impact of $s_{0}$   In Figure 5 (a), we tune $s_{0}$ with $k$ fixed to $20$. We first consider two extreme cases. In the first case, we set $s_{0}=0$, implying that $s<0.5$ all the time and SVD has dominant contribution in equation (5). As shown in “$s_{0}=0,k=20$” of Figure 5 (a), the dress still evenly sags down both sides of the body, unlike in the reference view where the dress flutters to the left side of the body. This indicates that without the sufficient guidance of geometry prior, the dynamics of each view will be independent of each other. Another extreme is when we set $s_{0}=1$, where SV3D has dominant contribution. In this case, it can no longer guarantee the consistency of details between different frames of invisible views. The comparison of the images of “$s_{0}=1,k=20$” shows a remarkable change in the shape of the tail of the hair. By further telescoping the results with different $s_{0}$, we finally adopt a compromised option to set $s_{0}$ to $0.5$, i.e., starting to drastically reduce the weight of the multi-view score at halfway of the denoising process. In most cases, this choice can achieve a satisfactory trade-off between smooth temporal transitions and geometrical consistency.
Impact of $k$   In addition, the decreasing speed of $s_{0}$ may also affect the quality of generated images, which is controlled by another parameter $k$. Therefore, we test the image generation under different $k$ given $s_{0}=0.5$ in Figure 5 (b). It can be seen that when $k$ is relatively small, the generated images may exhibit ghosting effects. On the other hand, when $k$ is too large, it may weaken the temporal smoothing effect brought by the video diffusion model in the early stage of denoising process. Consequently, we adopt $s=0.5,k=20$ as the default setting.
Impact of logistic schedule   Finally, we also examined the impact of different schedule types, as shown in Figure 5 (c). We compare with two counterparts: “Constant” ($s\equiv 0.5$) and “Linear” ($s=\frac{i}{N}$). It can be observed that other alternates would lead to ghosting artifacts.
5 Conclusion
In this work, we introduce Diffusion2, a novel framework for creating dynamic 3D content. This framework first generates a dense array of multi-view and multi-frame images with high parallelization, which are then used to build a full 4D representation through a reconstruction pipeline. The core assumption behind this framework is that the motion of an object viewed from one angle and its appearance from another are conditionally independent. Based on this assumption, we prove that we can directly sample synchronized multi-view videos in a denoising process by combining pretrained video and multi-view diffusion models. Our experimental results demonstrate the flexibility and effectiveness of our framework, showing that it can adapt to various prompts to produce high-quality 4D content efficiently and effectively. We hope that our work can inspire future research on unleashing and combining the geometrical and dynamic priors from foundation 3D and video diffusion models.
Appendix A Appendix
A.1 Discussions on the potential social impacts
In general, the proposed method has the potential to inspire a new approach to the automatic generation of dynamic 3D content. This could substantially lower the barriers to becoming an animator, save human resources within animation production, and ultimately reshape the industry’s landscape. Nonetheless, this transition may bring about issues such as unemployment and poverty for individuals whose livelihoods rely on these animation skills. Therefore, it may be necessary to gradually reduce the training scale in this industry and recalibrate the human resource supply to better align with the demands of future market.
On the other hand, the core contribution of Diffusion2 lies in reducing the 4D generation problem to the reconciliation of video and multi-view diffusion models. This indeed allows us to leverage recent advancements in these active fields effectively. However, the racial, gender, and cultural biases and stereotypes inherent in the foundational models may also be introduced into the generated 4D assets. Additionally, the potential abuse of generative models which may lead to copyright issues is also a matter of concern.
A.2 Limitations
As a preliminary exploration of zero-shot sampling of multi-view synchronized videos, this work still has some limitations. Overall, the performance of the proposed framework is severely constrained by the foundation generative models. For example, as can be seen from the second example in Figure 3 (a), the model has generated a weird textureless back for the cat. Additionally, due to biases in the training data of the video generation model, it is very unfriendly to thin structures in side views without clear semantics, easily generating unrealistic and inconsistent motions in these viewpoints. Furthermore, the motion generated by currently available video generation models may still be blurry or even corrupted. All these issues bring serious challenges to the subsequent reconstruction. However, we believe that the above problems will eventually be addressed with the further development of generative models, especially video generation models.
A.3 The applicability of the assumption and proof for the theorem in general case
@@
Another risk that needs to be pointed out is that our assumption about the conditional independence is too strong to apply in some cases. For example, if the target object rotates 180° over a period of time, the back view at the current moment and the front view after a period of time should not be conditionally independent given the front view at the current moment. However, existing video diffusion models typically generate videos from a single view without rotational movements, and our pipeline allows users to specify conditions to exclude such cases. Therefore, we have not discussed the potential limitations of this assumption in the main text for better readability and brevity of proof.
@@
It is worth noting that, even in such cases that violate assumption 3.1, our method can still work well. This is because the main theorem remains valid under these case and its correctness does not essentially rely on such a strong assumption. Actually, although extreme rotations might exist when considering frames with large intervals, they are unlikely to occur between adjacent frames in a natural video. Now considering that the distribution of $\hat{I}_{i,j}$ is almost entirely determined by its adjacent frames and views, we have the continuity assumption:
@@
Following the same derivation of the original proof 3.1, equations (6), (7) and (8) remain unchanged but all “$=$” in (7) and (8) should be $\approx$” because equation (7) is now established on equation (14) of assumption A.1 instead of assumption 3.1.
Then integrating new assumptions to correct equation (9), we obtain:
@@
This suggests the score of $I_{i,j}$ is only related to ^$\hat{\mathcal{I}}_{i_{\text{near}},j}$ and ^$\hat{\mathcal{I}}_{i,j_{\text{near}}}$. Similar to the derivation in equation (9), we can achieve:
@@
Now we have verified the correctness of the conclusion in equation 9. Based on the similar derivation, we can also ensure the safety of equation (7). This guarantees the correctness of our main theorem. ∎
A.4 Additional implementation details
@@
In the main text, we test the proposed approach on three types of prompts: reference image, single-view video, and multi-view images of static 3D model. The inputs used in our experiments for the first two tasks are respectively sourced from Animate124 zhao2023animate124 and Consistent4D jiang2023consistent4d , while that of the latter is collected from objaverse deitke2023objaverse . For the quantitative evaluation in the video-to-4D task, we follow Consistent4D’s testing protocol, reporting the average clip similarity across videos from four rendered views of all test examples. For the user study in the image-to-4D task, we collect the preferences of 50 participants on eight samples selected in Animate124 dataset, namely: astronaut-horse, cat-wave, eagle-fly, flag-wave, fox-game, lion-roar, monkey-bike, panda-dance.
Condition generation
Since there is a slight domain gap between the training data of video diffusion models and the background removed view-fixed single object video in our task, the video generation model within our framework is often responsible for most of the failure cases and corrupted areas. Inspired by FreeInit wu2023freeinit , we let informative spatial-temporal low-frequency components of the initial latent guide the generation for these out-of-domain videos to create the ideal layout in the early denoising stage. Specifically, we repeated the sampling process twice, and set the initial noise for the second sampling process as follows:
$x_{T}=\frac{x_{0}+\sigma_{T}\mathbf{n}}{\sqrt{1+\sigma_{T}^{2}}},$ (19)
where $x_{0}$ is the primary denoised latent, $\sigma_{T}$ is the initial noise level, and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the same Gaussian noise used to generate $x_{0}$ in the first sampling process. Experimental results indicate that this trick can effectively stabilize the generation, and the resulting significant reduction of failure cases cannot be achieved by trivially doubling the number of inference steps, which is aligned with the observations in wu2023freeinit . To minimize the influence on efficiency, we used only 25 steps in the first sampling process.
Image matrix generation
We also observed that the employed video diffusion model exhibits a noticeable view bias and struggles to stably generate side-view videos without clear semantics. Consequently, in the image matrix generation, we further adaptively tune the mixing weights for different views. Specifically, let $s_{i}$ and $r_{i}=1-s_{i}$ denote the weights of the video and multi-view score at the $i$-th sampling step in the logistic schedule. The weights adjusted by the view schedule are:
$s^{\prime}_{i}$ $=\alpha s_{i}$ (20)
$r^{\prime}_{i}$ $=1-\alpha+\alpha r_{i}$
$\alpha$ $=\frac{1}{2}\times\left(1+\cos\frac{2\pi\left(x+1\right)}{10.5}\right)$
where $v$ denotes the index of the view. This schedule reduces the contribution of the video generation model for the generation of side views.
A.5 Additional results
Figure 6: Text-to-4D results. We first generate an image conditioned on the text prompt and then perform our image-to-4D generation pipeline.
Text-to-4D generation
Naturally, Diffusion2 is also capable of text-to-4D generation. Although the text-to-video-to-4D approach sounds more reasonable since it allows control of the dynamics using natural language, existing video generation models fail to achieve satisfactory performance in practice. Therefore, for the text-to-4D task, we finally opted for the more stable text-to-image-to-4D generation approach, i.e., using image as the intermediary. Specifically, given a text prompt, we first synthesize the corresponding image. Subsequently, we perform background removal and recenter for the image to ensure the consistent use of the image extents with the training data of the multi-view diffusion model. Finally, the processed image is fed into our image-to-4D generation pipeline to synthesize the desired 4D assets. Actually, in Figure 3(a), there are already some input images that are AI-generated. In Figure 6, we demonstrate more generated 4D assets from text prompts.
4D reconstruction from partial views
To evaluate our method more comprehensively, we designed a quantitative benchmark, that assesses the fidelity of 4D reconstruction partial observation. Specifically, we render the front-view video and multi-view images at the reference frame from the animated 3D model as the condition for our generation pipeline. In Table 3, we report the PSNR and SSIM computed between generated frames and those rendered from six ground truth animated models collected from Maximo. For comparison, we evaluate the performance of Consistent4D under this setting. To adapt it to this task, we render the same view and timestamp using the camera intrinsic of Zero-1-to-3 liu2023zero , and incorporate pixel-wise reconstruction loss of reference views in the Consistent4D pipeline. From the qualitative comparisons in Figure 7, it can be seen that the baseline methods are inferior in terms of detail fidelity and geometric consistency with reference view. Moreover, we also demonstrate an advantage in the quantitative metrics presented in Table 3, which can boil down to a better alignment with ground truth camera parameters, as the qualitative metrics are largely dominated by the layout of the synthesized views in this task. These results once again confirm our superiority over the baseline method.
Figure 7: Partial reconstruction results. The synthesized images are conditioned on the reference single-view video and multi-view images.
Table 3: Quantitative comparison on the 4D reconstruction from partial observation.
[…]
Consistent4D jiang2023consistent4d 18.41 0.87
Diffusion2-stage1 20.45 0.89
Diffusion2-stage2 21.55 0.90

### v4

@@ A.2 Pseudocode for joint denoising of multi-view multi-frame image matrix via score composition
A.4 The applicability of the assumption 3.1 and proof for the theorem 3.1 in general case
@@
Another risk that needs to be pointed out is that our assumption about the conditional independence is too strong to apply in some cases. For example, if the target object rotates 180° over a period of time, the back view at the current moment and the front view after a period of time should not be conditionally independent given the front view at the current moment. Fortunately, existing video diffusion models typically generate videos from a single view without rotational movements, and our pipeline allows users to specify conditions to exclude such cases.
@@
In addition, it is worth noting that even in such cases that violate assumption 3.1, our method can still work well. This is because the main theorem remains valid under these case and its correctness does not essentially rely on such a strong assumption. Although extreme rotations might exist when considering frames with large intervals, they are unlikely to occur between adjacent frames in a natural video. Now considering that the distribution of $\hat{I}_{i,j}$ is almost entirely determined by its adjacent frames and views, we have the continuity assumption:
@@
Following the same derivation of the original proof A.3, equations (9), (10) and (11) remain unchanged but all “$=$” in (10) and (11) should be $\approx$” because equation 10 is now established on equation 15 of assumption A.1 instead of assumption 3.1.
Then integrating new assumptions to correct equation 12, we obtain:
@@
This suggests the score of $I_{i,j}$ is only related to ^$\hat{\mathcal{I}}_{i_{\text{near}},j}$ and ^$\hat{\mathcal{I}}_{i,j_{\text{near}}}$. Similar to the derivation in equation 12, we can achieve:
@@
Now we have verified the correctness of the conclusion in equation 12. Based on the similar derivation, we can also ensure the safety of equation 10. This guarantees the correctness of our main theorem. ∎
A.5 Additional experimental details
@@
In the main text, we test the proposed approach on two types of conditions: reference image and single-view video. The input images and videos used in the qualitative comparisons are sourced from the previous works Zhao et al. (2023); Jiang et al. (2024b); Tang et al. (2024b). For the quantitative evaluation in the video-to-4D task, we evaluate all methods on the Consistent4D test split and report the CLIP similarity, LPIPS, and FVD averaged on five views with accessible ground truth within the first 25 frames. Note that in equation 5, we introduced a coefficient $s$ to modulate the contribution of both models within the convex combination.
Scale schedule
Note that in equation 5, we introduced a coefficient $s$ to modulate the contribution of both models within the convex combination. Its schedule in denoising process can provide additional design space for enhancing the details at specific views. However, to ease the reproduction and reduce the reliance on manual priors, we fix it at 0.5 in all steps which is sufficient to achieve satisfactory results for most cases.
User study
Figure 7: The layout of our user study..
Figure 7 shows the template we used for user study. The participants are asked to pick one generation results from five candidates (displayed in video format) according to four dimensions: consistency with the reference video, multi-view consistency, temporal smoothness and overall model quality. Total 50 participants from diverse backgrounds are asked to do the questionnaire. And there are 11 cases for user study, including anya, cat-wave, patrick-star, flying-ironman, luigi, panda-dance, running-triceratops, sighing-frog, tiger-guitar, trump, walking-astronaut. All of them are collected from the previous works (Jiang et al., 2024b; Zhao et al., 2023; Tang et al., 2024b).
A.6 Additional results
Figure 8: Qualitative comparisons. For each method, we show images from two views and two timestamps.
Figure 9: More visualization results.
Figure 10: More visualization results.
In figure 8, we provede more cases for qualitative comparison. Similar conclusion to section 4.2 can be achieved that our results outperform the others. Figures 9 and 10 further shows additional results on custom data, with RGB and depth images from two views and three timestamps displayed. We recommend readers to watch the videos in our supplementary materials.